OpenStudy (anonymous):

Which of the following equations represents a line parallel to y=3/5x - 9 ?

OpenStudy (anonymous):

A. y=x/5-9 B. y= -5/3x+5 C. y= -3/5x - 9 D. y= 3/5x -6

hartnn (hartnn):

parallel lines have same slope! so, your equation will also be of the form y=3/5x+c

OpenStudy (anonymous):

AN equation of a line is said to be parrallel to another line when its slope is equal to each other

OpenStudy (anonymous):

Thanks. is it y=3/5x -6

hartnn (hartnn):

yes.

OpenStudy (anonymous):

as he said ^